OpenStudy (anonymous):

A baseball team won 10 games and lost 7 games. Approximately what percent of games did the team win?

OpenStudy (agent47):

10/17

OpenStudy (agent47):

59%

OpenStudy (agent47):

10/17 * 100%

OpenStudy (anonymous):

total games played =10+7=17 percentage of games won = 10/17*100 ans=58.8% estimated=59%